OpenStudy (anonymous):

Between what two consecutive integers is square root of 151? (A)12 and 13 (B)14 and 15 (C)11 and 12 (D)9 and 10

OpenStudy (anonymous):

9^ = 81 10^2 = 100 11^2 = 121 12^2 = 144 13^2 = 169

OpenStudy (anonymous):

its c @electrokid

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

oooo...its a lol